OpenStudy (anonymous):

how do you change -5 3/4 to an improper fraction?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

multiply the denominator by the whole number then add that to the numerator

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

so is the answer -17/4

OpenStudy (anonymous):

if it is, then yes

OpenStudy (anonymous):

moot?

OpenStudy (anonymous):

yeah?

OpenStudy (anonymous):

so is the answer -17/4?

OpenStudy (anonymous):

thats right good job

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

anytime :)

OpenStudy (asnaseer):

the answer is not right

OpenStudy (asnaseer):

\(-5\frac{3}{4}=-(5+\frac{3}{4})\) what you have calculated is: \(-5+\frac{3}{4}\)